April 5, 2018 Drift Chamber meeting
- Instructions for Bluejeans meeting connection
- Meeting ID: 290664653
- CDC update
- FDC update
Participants: Sean(FSU), Eugene, Beni, Simon, Alex A., Alex B., Sergey, and Lubomir (JLab).
- Sergey showed an updated version of his report about the CDC background studies. It includes plots of the hit occupancy as function of the beam current with physics trigger and pulser (random) for inner and outer rings. While the two curves (physics and random) are parallel for the outer rings, at high occupancy as it is in the inner ring, this is not the case. There's no quantitative description yet. The plan is to take more data at higher intensities.
- The phi-asymmetry in the inner CDC rings that Sergey found in case of no radiator and that is significantly suppressed when inserting W-foil (40 um) or emptying the target, can be explained by a synchrotron radiation from the vertically-bending magnet in the ramp to the tagger (Richard proposed this).
- Naomi suggested looking in the phi-asymmetry of the pedestals, that may be coming from the same polarized synchrotron radiation.
- Similar noise that has been observed last spring appeared again on two FDC wire cards (connected to one TDC module). The noise is present only after starting the DAQ and disappears soon after you stop it. Beni suggested disconnecting both cables from the TDC when looking with a scope on one of them. This was done later, but no difference in the noise was found.